The management of Quest Media Inc. is considering two capital investment projects. The estimated net cash flows from each project are as follows:

Year Radio Station TV Station
1 $420,000 $880,000
2 420,000 880,000
3 420,000 880,000
4 420,000 880,000
Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

The radio station requires an investment of $1,199,100, while the TV station requires an investment of $2,278,320. No residual value is expected from either project.

Required:

1a. Compute the net present value for each project. Use a rate of 10% and the present value of an annuity of $1 in the table above. If required, use the minus sign to indicate a negative net present value. If required, round to the nearest whole dollar.

Radio Station TV Station
Present value of annual net cash flows $ $
Less amount to be invested $ $
Net present value $ $

1b. Compute a present value index for each project. If required, round your answers to two decimal places.

Present Value Index
Radio Station
TV Station

2. Determine the internal rate of return for each project by (a) computing a present value factor for an annuity of $1 and (b) using the present value of an annuity of $1 in the table above. If required, round your present value factor answers to three decimal places and internal rate of return to the nearest whole percent.

Radio Station TV Station
Present value factor for an annuity of $1
Internal rate of return % %
Answer #1
1a
Radio Station TV Station
Present value of annual net cash flows 1331400 2789600
Less amount to be invested 1199100 2278320
Net present value 132300 511280
b
Present Value Index
Radio Station 1.11 =1331400/1199100
TV Station 1.22 =2789600/2278320
2
Radio Station TV Station
Present value factor for an annuity of $1 2.855 2.589
Internal rate of return 15% 20%
Formulas used:
Radio Station TV Station
Present value of annual net cash flows =420000*3.170 =880000*3.170
Radio Station TV Station
Present value factor for an annuity of $1 =1199100/420000 =2278320/880000